# Do You Hear What I Mean? Quantifying the Instruction-Perception Gap in Instruction-Guided Expressive Text-To-Speech Systems

### v2

@@ 1 Introduction
Instruction-guided text-to-speech (ITTS) [1, 2] enables users to steer speech synthesis using natural-language prompts (e.g., “read this joyfully,” “speak like a child”). This approach offers a transparent and flexible alternative to conventional TTS pipelines [3, 4] that often require low-level acoustic controls or specialized labels for prosody and timing. By shifting control to free-form language, ITTS promises to enhance accessibility for content creation, assistive technologies, education, and interactive media.
@@
To address this question, we introduce a novel evaluation of ITTS controllability. Our study is the first to incorporate adverbs of degree (e.g., slightly, extremely) and graded emotion intensity (e.g., ecstatic, happy) as explicit evaluation dimensions. We also present the first large-scale collection of human perceptual ratings for speaker age and word-level emphasis. To systematically examine the gap between instructions and listener perception across these dimensions, we developed a new analysis framework and compiled the Expressive VOice Control (E-VOC) corpus, consisting of high-quality judgments from over 165 human raters. To ensure the reliability of our findings and enable reproducibility, all data were gathered through a quality-controlled process, and we will publicly release both the corpus and the analysis pipeline.
@@ 2.1 ITTS Systems and Selection
The field of Instruction-guided Text-to-Speech (ITTS) has seen rapid advancement, with many models capable of generating speech from descriptive prompts [1]. Although robust systems such as Audiobox [9] exist, their closed-source nature limits the transparency and reproducibility necessary for this study. Therefore, to ensure a comprehensive and replicable analysis, we selected five representative models across three distinct categories. First, to represent the state-of-the-art in open-source research, we included Parler-TTS [10] and PromptTTS++ [11]. These models are publicly available, allowing for the in-depth analysis required for this study. Second, to represent the leading edge of commercial ITTS systems, we incorporate GPT-4o-mini-TTS [12]. Its efficient and high-quality API provides an insightful analysis for production-grade expressive synthesis. Finally, to test the capabilities of non-specialized systems, we included UniAudio [13], a unified audio generative model. Its inclusion allows us to assess whether a general-purpose audio foundation model can achieve perceptual alignment comparable to others.
@@ 3.1.1 Proposed Control Dimensions
Task I. Adverbs of Degree (Adv. Deg.) tests whether models follow degree modifiers such as “slightly,” “very,” and “extremely” to adjust prosody (loudness, pitch, speaking rate). This dimension is important because adverbial scaling provides users with a simple way to control the fine-grained prosodic variation, essential for storytelling and emotional expression.
@@ 3.1.2 Other Control Dimensions
extremelyverymoderateslightly1.51.5222.52.533Intensity LevelAngryextremelyverymoderateslightlyHappyextremelyverymoderateslightlySurprisedextremelyverymoderateslightlySadUpsetFrustratedIrritatedAngryOutraged1.51.5222.52.533Intensity LevelSatisfiedContentHappyOverjoyedEcstaticIntriguedUncertainAmazedStunnedSurprisedGloomyDisappointedUnhappySadHeartbroken
@@ 4.2 Emotion–Intensity Adjectives
Other models show weaker or flatter distinctions. Parler-large and Parler-mini captured some variation, but the perceptual steps between adjacent adjectives were small. PromptTTS++ often produced nearly indistinguishable outputs across terms, while UniAudio occasionally exhibited reversed trends, with listeners rating mid-level adjectives as more intense than stronger ones. These results suggest that while most ITTS systems can generate distinct categorical emotions, only gpt-4o can reliably control their fine-grained emotion intensity.
@@ 4.4 Word-level Emphasis
Controlling word-level emphasis was a significant challenge for five ITTS systems. As shown in Table 3, gpt-4o achieved the highest accuracy, yet its score of 0.265 indicates that even the best-performing model struggled. This task highlights a critical area for improvement, as effective emphasis requires precise and consistent coordination of pitch excursion, duration, and intensity at the word level.
@@ 5 Conclusion and Future Work
Conclusion. This work addresses the largely unexplored relationship between natural-language instructions and listener perception in advanced ITTS systems. We introduced a novel evaluation framework centered on graded emotion intensity, using both adverbs of degree (e.g., “slightly happy”) and emotionally ordered adjectives (e.g., from “Content” to “Happy” to “Ecstatic”) to measure fine-grained control. Our empirical analysis of five leading ITTS models revealed two key patterns: (1) Among the systems tested, gpt-4o was the only one that reliably translated degree modifiers and graded adjectives into perceptually ordered changes in loudness, pitch, speaking rate, and emotion intensity. (2) Fine-grained controls like word-level emphasis and speaker age were inconsistently realized across all models. Most used ITTS systems defaulted to adult-like voices and produced weak or unstable emphasis cues. To sum up, while contemporary ITTS can follow some high-level styles with coarse reliability, achieving consistent and fine-grained control that aligns with human perception remains a significant open challenge.

### v3

@@ 1 Introduction
Instruction-guided text-to-speech (ITTS) [1, 2] enables users to steer speech synthesis using natural-language prompts (e.g., “read this joyfully” or “speak like a child”). This approach offers a transparent and flexible alternative to conventional TTS pipelines [3, 4] that often require low-level acoustic controls or specialized labels for prosody and timing. By shifting control to free-form language, ITTS promises to enhance accessibility for content creation, assistive technologies, education, and interactive media.
@@
To address this question, we introduce a novel evaluation of ITTS controllability. Our study is the first to incorporate adverbs of degree (e.g., slightly, extremely) and graded emotion intensity (e.g., ecstatic, happy) as explicit evaluation dimensions. We also present the first large-scale collection of human perceptual ratings for speaker age and word-level emphasis. To systematically examine the gap between instructions and listener perception across these dimensions, we developed a new analysis framework and compiled the Expressive VOice Control (E-VOC) corpus [^1] Project Website, consisting of high-quality judgments from over 165 human raters. To ensure the reliability of our findings and enable reproducibility, all data were gathered through a quality-controlled process, and we will publicly release both the corpus and the analysis pipeline.
@@ 2.1 ITTS Systems and Selection
The field of Instruction-guided Text-to-Speech (ITTS) has seen rapid advancement, with many models capable of generating speech from descriptive prompts [1]. Although robust systems such as Audiobox [9] exist, their closed-source nature limits transparency and reproducibility, which are essential for this study. Therefore, to ensure a comprehensive and replicable analysis, we selected five representative models across three distinct categories. First, to represent the state-of-the-art in open-source research, we included Parler-TTS [10] and PromptTTS++ [11]. These models are publicly available, allowing for the in-depth analysis required for this study. Second, to represent the leading edge of commercial ITTS systems, we incorporate GPT-4o-mini-TTS [12]. Its efficient and high-quality API provides an insightful analysis for production-grade expressive synthesis. Finally, to test the capabilities of non-specialized systems, we included UniAudio [13], a unified audio generative model. Its inclusion allows us to assess whether a general-purpose audio foundation model can achieve perceptual alignment comparable to others.
@@ 3.1.1 Proposed Control Dimensions
Task I. Adverbs of Degree (Adv. Deg.) tests whether models follow degree modifiers such as “slightly,” “very,” and “extremely” to adjust prosody (loudness, pitch, speaking rate) and emotion. This dimension is important because adverbial scaling provides users with a simple way to control the fine-grained prosodic variation, essential for storytelling and emotional expression.
@@ 3.1.2 Other Control Dimensions
extremelyveryN/Aslightly1.51.5222.52.533Intensity LevelAngryextremelyveryN/AslightlyHappyextremelyveryN/AslightlySurprisedextremelyveryN/AslightlySadUpsetFrustratedIrritatedAngryOutraged1.51.5222.52.533Intensity LevelSatisfiedContentHappyOverjoyedEcstaticIntriguedUnexpectedAmazedStunnedSurprisedGloomyDisappointedUnhappySadHeartbroken
@@ 4.2 Emotion–Intensity Adjectives
Other models show weaker distinctions. Parler-large and Parler-mini captured some variation, but the perceptual steps between adjacent adjectives were small. PromptTTS++ often produced nearly indistinguishable outputs across terms, while UniAudio occasionally exhibited reversed trends, with listeners rating mid-level adjectives as more intense than stronger ones. Overall, most ITTS systems can generate distinct categorical emotions, but only gpt-4o reliably controls fine-grained emotion intensity.
@@ 4.4 Word-level Emphasis
Controlling word-level emphasis was a significant challenge for five ITTS systems. As shown in Table 3, gpt-4o achieved the highest accuracy, yet its score of 0.265 indicates that even the best-performing model struggled. Since each sentence contained 6 candidate words on average, random guessing would yield an expected accuracy of only $1/7\approx 0.143$, including the option ”no word emphasized”. This task highlights a critical area for improvement, as effective emphasis requires precise and consistent coordination of pitch excursion, duration, and intensity at the word level.
@@ 5 Conclusion and Future Work
Conclusion. This work addresses the largely unexplored link between natural-language instructions and listener perception in ITTS. We proposed a novel framework for evaluating fine-grained control using adverbs of degree (e.g., “slightly happy”) and ordered emotional adjectives (e.g., from “Content” to “Happy” to “Ecstatic”). Analysis of five leading ITTS models revealed two key patterns: (1) Among the systems tested, gpt-4o was the only one that reliably translated degree modifiers and graded adjectives into perceptually ordered changes in loudness, pitch, speaking rate, and emotion intensity. (2) Fine-grained controls like word-level emphasis and speaker age were inconsistently realized across all models. Most ITTS systems defaulted to adult-like voices and produced weak emphasis cues. To sum up, while current ITTS can follow some high-level styles with coarse reliability, achieving consistent and fine-grained alignment with human perception remains an open challenge.